Arshadi Corp.'s sales last year were $67,000, and its total assets were $22,000. What was its total assets turnover ratio (TATO)?

Select the correct answer.

Arshadi Corp.'s sales last year were $67,000, and its total assets were $22,000. What was its total assets turnover ratio (TATO)?

Select the correct answer.

a. 1.85

b. 3.05

c. 3.45

d. 2.25

e. 2.65

Orono Corp.'s sales last year were $585,000, its operating costs were $362,500, and its interest charges were $12,500. What was the firm's times interest earned (TIE) ratio?

Select the correct answer.

a. 15.80

b. 18.80

c. 14.80

d. 17.80

e. 16.80

Branch Corp.'s total assets at the end of last year were $350,000 and its net income after taxes was $22,750. What was its return on total assets?

Select the correct answer.

a. 6.80%

b. 7.10%

c. 6.50%

d. 6.20%

e. 5.90%

Nikko Corp.'s total common equity at the end of last year was $325,000 and its net income after taxes was $60,000. What was its ROE?

Select the correct answer.

a. 19.96%

b. 17.96%

c. 19.46%

d. 18.46%

e. 18.96%

Vang Corp.'s stock price at the end of last year was $46 and its earnings per share for the year were $2.30. What was its P/E ratio?

Select the correct answer.

a. 19.20

b. 20.80

c. 19.60

d. 20.00

e. 20.40

An investor is considering starting a new business. The company would require $625,000 of assets, and it would be financed entirely with common stock. The investor will go forward only if she thinks the firm can provide a 13.5% return on the invested capital, which means that the firm must have an ROE of 13.5%. How much net income must be expected to warrant starting the business?

Select the correct answer.

a. $84,274

b. $84,375

c. $84,325

d. $84,224

e. $84,174
Answer #1
Total assets turnover ratio = Net sales / Total assets
Total assets turnover ratio = 67000/22000 i.e 3.05
Times interest earned ratio = Net income before interest and tax / Interest expense
Times interest earned ratio = 221500/12500 i.e 17.80
Return on total assets = Net income after taxes / total assets
Return on total assets = 22750/ 350000 i.e 0.065 or 6.50%
Return on equity = Net income after taxes / Common equity
Return on equity = 60000/325000 i.e 0.1846 or 18.46%
Price earning ratio = Price / Earning
Price earning ratio = 46 / 2.30 i.e 20
Return on equity = Return / Total equity
13.5% = Net income / Total equity
13.5% = Net Income / 625000
Net income = 625000*13.5% i.e 84375